# Gravitational wave probes of particle dark matter: a review

### v3

@@ II.5 Tensor dark matter
Modifications to gravity due to an additional spin-2 particle could act as DM. Specifically, bimetric gravity Hassan and Rosen (2012); Schmidt-May and von Strauss (2016), a theory in which a massless and massive spin-2 field interact, provides a plausible candidate for ultralight dark matter Marzola et al. (2018), produced by the misalignment mechanism Preskill et al. (1983).
Here, we consider a massive spin-2 field $M_{\mu\nu}$ described by the Fierz-Pauli Lagrangian density Armaleo et al. (2020):
@@
For the Friedman-Lemaitre-Robertson-Walker (FLRW) background metric, the equations of motion for this field, assuming it is ultralight, can be derived at late times as in  Marzola et al. (2018); Aoki and Maeda (2018); Kun et al. (2018). The field spatial component can be written in a similar way as for dark photons and scalar bosons at given position $\mathbf{x}$:
@@
$M_{ij}$ $=\frac{\sqrt{2\rho_{\text{DM}}(\vec{x})}}{m}\cos{\left(\frac{mc^{2}}{\hbar}t+\vec{k}\cdot\vec{x}+\Upsilon(\vec{x})\right)}\varepsilon_{ij}(\vec{x})\,,$ (58)
where $\Upsilon$ is a random phase, and $\varepsilon_{ij}$ encodes information about the five polarizations of the massive spin-2 field.
@@
The strain on the GW detectors arises analogously to that from GWs: a stretching of space-time in the presence of the field, since the massive spin-2 metric and its coupling constant can be absorbed into the definition of the massless spin-2 metric in the linear regime of the coupling constant $\alpha_{\rm TB}$ [^4] In the literature, $\alpha_{\rm TB}=\alpha$ simply, but to disentangle it from the fine structure constants, we rename it. TB is tensor boson. Armaleo et al. (2020); thus, the perturbation $h_{ij}$ is:
@@
$h_{ij}(t)$ $=\frac{\alpha_{\rm TB}}{M_{\text{Pl}}}M_{ij}(t)=\frac{\alpha_{\rm TB}\sqrt{2\rho_{\text{DM}}}}{mM_{\text{Pl}}}\cos{\left(\frac{DMc^{2}}{\hbar}t+\Upsilon\right)}\varepsilon_{ij}(\mathbf{x})\,,$ (59)
@@
where $\alpha$ is a dimensionless constant that quantifies the difference between the strengths of each of the spin-2 fields, and $M_{\text{Pl}}$ is the Planck mass.
Contracting $h_{ij}$ with the GW detector response function $D^{ij}$ in the tensor case, the resulting strain is:
@@
$h(t)=D^{ij}h_{ij}=\frac{\alpha_{\rm TB}\sqrt{\rho_{\text{DM}}}}{\sqrt{2}mM_{\text{Pl}}}\cos{\left(\frac{m_{\rm DM}c^{2}}{\hbar}t+\Upsilon\right)}\Delta\varepsilon,$ (60)
where $\Delta\varepsilon:=\,\varepsilon_{ij}(n^{i}n^{j}-m^{i}m^{j})$.
II.6 Gravitational interactions of dark matter
@@
The previous sections assumed a coupling of DM to standard-model particles. However, in principle DM need not interact at all with the standard model; thus, models of DM have been devised in which only a minimal coupling to gravity is assumed. This is, in fact, the simplest assumption that can be made, since we only have evidence for the existence of DM through gravitational interactions.
One such hypothetical scalar DM particle has a Lagrangian of:
@@
and Khmelnitsky and Rubakov (2014):
@@
where $\gamma(\vec{x})$ is a phase that depends on position $\vec{x}$, and $\hat{\phi}(\vec{x})$ describes the interference pattern caused by the ultralight dark matter. As before, the scalar field density is normalized by $\rho_{\text{DM}}$.
Scalar DM could cause the arrival times of pulses from millisecond pulsars on earth to change, in an analogous way that GWs do Khmelnitsky and Rubakov (2014). Such a difference in the arrival times can be written as Khmelnitsky and Rubakov (2014); Porayko et al. (2018):
@@
where $\gamma_{\text{P}}\equiv 2\gamma(\vec{x}_{\text{p}})-2m_{\rm DM}d_{p}/c$ ($\gamma_{\text{E}}\equiv 2\gamma(\vec{x}_{\text{e}})$) parameterize random phases evaluated at the pulsar (P) or Earth (E), and $d_{p}$ is the distance between the pulsar and the Earth. The DM energy density background is assumed to be constant when calculating $\Psi_{\text{c}}(\vec{x})$, though the possibility of deviations, due to the coherently oscilalting ultralight dark-matter field, are parametrized in terms of two phase factors: one for the pulsar $\hat{\phi}^{2}(\vec{x}_{\text{p}})\equiv\hat{\phi}^{2}_{\text{P}}$ and one for the Earth ($\hat{\phi}^{2}(\vec{x}_{\text{e}})\equiv\hat{\phi}^{2}_{\text{E}}$). Since all the pulsars are separated by at least a kpc, assuming a constant DM energy density is valid Porayko et al. (2018).
III Observational constraints on dark matter directly interacting with gravitational-wave interferometers
Section II focused on the currently-known types of DM that could cause a signal in GW interferometer data. While we have indicated the strain amplitudes of such signals for some of the DM models considered, we have not yet discussed whether the interferometers are actually sensitive to such weak signals. In this section, we will describe how searches for DM interactions with GW detectors are performed, show current constraints on various models, and explain what future constraints will be on others.
Each method discussed below can achieve a certain sensitivity, that is, a minimum detectable strain amplitude on the detector at a given confidence level. This strain amplitude, the left-hand side of many equations in Section II, is then mapped to constraints on the physical coupling constant of DM to the standard model. The methods themselves are agnostic towards the particular model of DM that is being searched for.
@@ III.1 Methods
If the signal were purely monochromatic, we could simply take a single fast Fourier transform of the data and look for peaks in the power spectrum. However, the stochastic frequency variations prevent us from doing that. If we observe for a duration $T_{\text{obs}}$ longer than $T_{\text{coh}}$, the signal will not be sinusoidal and thus its power will be spread among different frequency bins, which would inhibit a possible detection. Thus, the following methods have been designed to combine signal power across chunks of data $T_{\text{FFT}}\ll T_{\text{obs}}$ in a way that optimizes sensitivity towards particle DM interactions with the standard model and gravity.
@@ III.1.1 Cross-correlation
Cross-correlation Pierce et al. (2018); Guo et al. (2019) requires at least two separate time-domain datasets that are Fourier transformed and multiplied together to compute the cross-power in each frequency bin. The cross power is then divided by the auto-power of each detector, and then summed over all the Fourier transforms to arrive at a measure of power at each frequency. It is then divided by the standard deviation of the noise to compute the signal-to-noise ratio.
The cross-correlated signal strength for detector pair $IJ$ is:
@@
in the $j^{\rm th}$ frequency bin at the $i^{\rm th}$ time. “*” denotes the complex conjugate, and $z_{1,ij}$ and $z_{2,ij}$ is the Fourier transform of the time-domain data from detectors 1 and 2, respectively, and $N_{\text{FFT}}$ is the total number of FFTs taken over the observing run. If the observing run lasts $T_{\text{obs}}$ and the FFT length is $T_{\text{FFT}}$, then $N_{\text{FFT}}=[T_{\text{obs}}/T_{\text{FFT}}]$. $P_{1,ij}$ and $P_{2,ij}$ are the individual detector power spectral densities, which are typically estimated using a running median of each FFT. Typically $T_{\text{FFT}}=1800$ s, independently of the DM mass, which ensures a relatively easy analysis but implies the sensitivity towards DM interactions is not optimal at each mass analyzed.
To obtain a detection statistic, i.e. the signal-to-noise ratio, we have to divide the signal strength by the variance of the data, which is:
@@
The signal-to-noise ratio (SNR), the dection statistic, is then simply:
@@
In the presence of pure Gaussian noise, the SNR will follow a normal distribution, with a mean of 0 and a standard deviation of 1. If the SNR exceeds a certain threshold, which is set both theoretically and by the trials factor, then a particular frequency is classified as being “significant”.
Depending on the type of DM searched for, the signal-to-noise ratio for spin-0 (labeled SD), spin-1 (DP) and spin-2 (DG) cross-correlation searches can be written as Manita et al. (2024):
${\rm SNR}^{\rm SD}_{IJ,j}$ $\approx\left|\frac{2\sqrt{2}\alpha_{A}^{2}f_{\rm SD}T_{\rm eff}({\cal B}^{\prime}_{IJ}\Gamma_{IJ}+{\cal C}^{\prime}\gamma_{IJ})}{\sqrt{S_{n,I}(\frac{m_{\rm DM}}{2\pi})S_{n,J}(\frac{m_{\rm DM}}{2\pi})}}\right|\,.$ (68)
${\rm SNR}^{\rm DP}_{IJ,j}$ $\approx\left|\frac{2\sqrt{2}\epsilon_{D}^{2}f_{\rm DP}T_{\rm eff}({\cal B}_{IJ}\Gamma_{IJ}+{\cal C}(1+\frac{1}{3}\Omega^{z}_{\rm DP})\gamma_{IJ})}{\sqrt{S_{n,I}(\frac{m_{\rm DM}}{2\pi})S_{n,J}(\frac{m_{\rm DM}}{2\pi})}}\right|\,,$ (69)
${\rm SNR}^{\rm DG}_{IJ,j}$ $\approx\left|\frac{2\sqrt{2}\alpha^{2}f_{\rm DG}T_{\rm eff}{\cal A}\gamma_{IJ}}{\sqrt{S_{n,I}(\frac{m_{\rm DM}}{2\pi})S_{n,J}(\frac{m_{\rm DM}}{2\pi})}}\right|\,,$ (70)
For spin-0: $f_{\rm SD}\equiv\rho_{\rm SD}/\rho_{\text{DM}}$ is the ratio of the energy density of the spin-0 DM to the total energy density of DM. $\alpha_{A}$ is the dimensionless coupling constant that depends on the coupling constants of the scalar DM particles to standard model ones, e.g. the electron and photon, and are given by Eq. 44-47 in Manita et al. (2024). $\gamma_{IJ}$ is the overlap-reduction function that quantities the sensitivity loss by cross correlating detectors that are not co-located and co-aligned, and is given in Tab. I in Manita et al. (2024) for different pairs of detectors. It specifically arises from the spatial displacement of mirrors by DM. $\Gamma_{IJ}$ is the overlap-reduction function that arises from the finite-light time travel effects of the signal on the mirrors Morisaki et al. (2021).
For spin-1: $f_{\rm DP}\equiv\rho_{\rm DP}/\rho_{\text{DM}}$ is the ratio of the energy density of the spin-1 DM to the total energy density of DM. $\Omega^{z}_{\rm DP}$ is a term that depends on the polarization of DM, and $\Omega^{z}_{\rm DP}=1/3$ is when the polarizations are randomized.
For spin-2: The SNR is independent of how the DM energy density is distributed in each polarization mode. $f_{\rm DG}\equiv\rho_{\rm DG}/\rho_{\text{DM}}$ is the ratio of the energy density of the spin-2 DM to the total energy density of DM.
In all the expressions $T_{\rm eff}=T_{\text{obs}}$ if $T_{\text{obs}}<T_{\text{coh}}$, and $T_{\rm eff}=\sqrt{T_{\text{obs}}T_{\text{FFT}}}$ if $T_{\text{obs}}>T_{\text{coh}}$. Furthermore, the definitions of the dimensionless coefficients ${\cal C}$, ${\cal B}_{IJ}$ and ${\cal A}$ are:
${\cal C}$ $\equiv\frac{\rho_{\rm DM}e^{2}v_{0}^{2}}{5\omega^{2}}\left(\frac{Q_{D}}{M}\right)^{2}\,,$ (71)
${\cal B}_{IJ}$ $\equiv\frac{8\rho_{\rm DM}e^{2}}{3\omega^{4}L_{I}L_{J}}\left(\frac{Q_{D}}{M}\right)^{2}\sin^{2}\left(\frac{\omega L_{I}}{2}\right)\sin^{2}\left(\frac{\omega L_{J}}{2}\right)\,,$ (72)
${\cal A}$ $\equiv\frac{\rho_{\rm DM}}{5\omega^{2}M_{\text{Pl}}^{2}}$ (73)
where $\omega=2\pi f$ is the angular frequency of the ultralight dark matter, $L_{I}$ and $L_{J}$ are the arm-lengths of each interferometer,
III.1.2 Excess power time-frequency method

### v4

@@ II.5 Tensor dark matter
Modifications to gravity due to an additional spin-2 particle could act as DM. Specifically, bimetric gravity Hassan and Rosen (2012); Schmidt-May and von Strauss (2016), a theory in which a massless and massive spin-2 field interact, provides a plausible candidate for ultralight dark matter Marzola et al. (2018).
Here, we consider a massive spin-2 field $M_{\mu\nu}$ described by the Fierz-Pauli Lagrangian density Fierz and Pauli (1939); Armaleo et al. (2020):
@@
For the Friedman-Lemaitre-Robertson-Walker (FLRW) background metric, the equations of motion for this field, assuming it is ultralight, can be derived at late times as in  Marzola et al. (2018); Aoki and Maeda (2018); Kun et al. (2018). The field spatial component can be written in a similar way as for dark photons and scalar bosons at given position:
@@
$M_{ij}$ $=\frac{\sqrt{2\rho_{\text{DM}}(\vec{x})}}{m_{\rm DM}}\cos{\left(m_{\rm DM}t+\vec{k}\cdot\vec{x}+\Upsilon(\vec{x})\right)}\varepsilon_{ij}(\vec{x})\,$, (57)
where and $\varepsilon_{ij}$ encodes the polarizations of the massive spin-2 field.
@@
The strain on GW detectors arises analogously to that from GWs: a stretching of space-time in the presence of the field, since the massive spin-2 metric and its coupling constant can be absorbed into the definition of the massless spin-2 metric in the linear regime of the coupling constant $\alpha_{\rm TB}$ [^5] In the literature, $\alpha_{\rm TB}=\alpha$ simply, but to disentangle it from the fine structure constants, we rename it. TB is tensor boson. Armaleo et al. (2020); thus, the perturbation $h_{ij}$ is:
@@
$h_{ij}(t)$ $=\frac{\alpha_{\rm TB}}{M_{\text{Pl}}}M_{ij}(t)=\frac{\alpha_{\rm TB}\sqrt{2\rho_{\text{DM}}}}{m_{\rm DM}M_{\text{Pl}}}\cos{\left(m_{\rm DM}t+\Upsilon\right)}\varepsilon_{ij}(\vec{x})\,$, (58)
@@
where $\alpha_{\rm TB}$ is a dimensionless constant that quantifies the difference between the strengths of each of the spin-2 fields, and $M_{\text{Pl}}$ is the Planck mass.
To obtain the strain on GW interferometers, we note that the detector response function can be parameterized in terms of a tensor, often written as $D^{ij}$ Romano and Cornish (2017). By contracting $h_{ij}$ with $D^{ij}$, the resulting strain can be derived Armaleo et al. (2021).
@@
$h(t)=D^{ij}h_{ij}=\frac{\alpha_{\rm TB}\sqrt{\rho_{\text{DM}}}}{\sqrt{2}mM_{\text{Pl}}}\cos{\left(m_{\rm DM}t+\Upsilon\right)}\Delta\varepsilon,$ (59)
where $\Delta\varepsilon:=\,\varepsilon_{ij}(n^{i}n^{j}-m^{i}m^{j})$ and $n,m$ are unit vectors pointing along directions $i,j$.
We note that an interesting discrepancy arises when comparing the results from Refs. Aoki and Mukohyama (2016); Manita et al. (2024) with those from Refs. Marzola et al. (2018); Armaleo et al. (2021) regarding the strain induced by tensor dark matter. The two sets of works differ cause a factor of 2 difference in the strain, which can be traced back to a factor of eight difference in their respective Lagrangians. This difference stems from variations in the normalization factors and scaling of the tensor field amplitude, as well as differences in how the tensor field’s energy density is treated. Such discrepancies are not uncommon in the study of dark matter and may have implications for the interpretation of experimental constraints and the modeling of astrophysical signals. However, we note that no constraints on tensor DM have been placed as of yet using GW interferometers.
II.6 Gravitationally interacting dark matter
@@
The previous sections assumed a coupling of DM to standard-model particles. However, in principle DM need not interact at all with the standard model; thus, models of DM have been devised in which only a minimal coupling to gravity is assumed. This is, in fact, the simplest assumption that can be made, since we only have evidence for the existence of DM through gravitational interactions. Below, we consider two kinds of scalar DM that couple only to gravity.
II.6.1 Minimally coupled dark matter
In minimal coupling models, DM interacts with the standard model exclusively through gravity by contributing to the stress-energy tensor. The gravitational field responds to this energy density and pressure according to Einstein’s field equations, where the presence of dark matter modifies the curvature of spacetime in the same manner as any other form of energy or matter. Here, ultralight dark matter can source time-varying gravitational potentials, leading to observable effects such as oscillations in pulsar timing residuals, which will be discussed in Section III.3.
A commonly studied example of a minimally coupled scalar DM field $\phi$ is described by the Lagrangian Khmelnitsky and Rubakov (2014):
@@ II.6.1 Minimally coupled dark matter
where $g^{\mu\nu}$ is the metric tensor, and $g$ is the determinant of the metric tensor. In this model, the scalar field evolves:
@@
where $\gamma(\vec{x})$ is a phase that depends on position $\vec{x}$, and $\hat{\phi}(\vec{x})$ describes the variation in space of the ultralight dark-matter field. As before, the scalar field density is normalized by $\rho_{\text{DM}}$.
Scalar DM can cause the arrival times of pulses from millisecond pulsars on earth to change, in an analogous way that GWs do Khmelnitsky and Rubakov (2014), i.e. introduce stochastic fluctuations of the metric Kim and Mitridate (2024); Kim (2023). Such a difference in the arrival times can be written as Khmelnitsky and Rubakov (2014); Porayko et al. (2018):
@@
[…]
II.6.2 Non-minimally coupled (conformal) dark matter
Another form of DM that couples only to gravity would be conformal DM. This kind of DM typically involve scalar fields that couple to the trace of the energy-momentum tensor. These models are closely related to the dilaton-type scalar DM discussed in Section II.3, but with specific couplings motivated by scalar–tensor theories of gravity, such as Fierz–Jordan–Brans–Dicke (FJBD) and Damour–Esposito–Farese (DEF) theories Fierz and Pauli (1939); Fierz (1956); Dicke (1962); Jordan (1959); Damour and Esposito-Farese (1992, 1993).
In such models, DM couples universally to the gravitational sector through a conformal rescaling of the metric. This coupling can induce a gravity-mediated force between neutron stars and the ultralight dark matter field, altering the moment of inertia of rotating pulsars – a manifestation of the Nordtvedt effect Nordtvedt (1968). These variations lead to deterministic changes in pulsar ToA and constitute a violation of the equivalence principle. This kind of conformal scalar DM has been extensively studied and constrained using binary pulsar timing measurements Damour and Taylor (1992); Kramer et al. (2006), which will be discussed inSection III.3.
In the case of a linear FJBD coupling, the pulsar timing residuals take the form:
$\Delta t(t)$ $=2\alpha s_{I}\left.\frac{\sqrt{\rho_{\text{DM}}}}{M_{\text{Pl}}m_{\rm DM}^{2}}\hat{\phi}(\vec{x})\sin(m_{\rm DM}t+\theta(\vec{x}))\right|_{t_{\text{start }}-d}^{t_{\text{end }}-d},$ (64)
where $s_{I}$ is the angular moment sensitivity parameter computed in Kuntz and Barausse (2024) and is of $\mathcal{O}(1)$ in FJBD theory, $\alpha$ is the strength of the coupling, $\theta(\vec{x})$ is a random phase, and $\hat{\phi}$ is a stochastic parameter whose value depends on whether the correlated, uncorrelated or pulsar correlated scenarios are considered (see Section III.3 for a description of these terms).
In summary, both dilaton-type and conformal scalar DM predict couplings that can modify either the properties of matter or the effective gravitational interaction, leading to observable signatures. In ground-based interferometers, these effects manifest primarily as strain signals induced by oscillations in material properties or cavity frequencies, while in pulsar timing arrays they appear as deterministic modulations in pulse arrival times due to variations in neutron-star moment of inertias or equivalence-principle–violating forces. These complementary channels allow terrestrial and astrophysical detectors to probe overlapping but distinct regions of parameter space. In the next section, we review the current observational constraints on such models, focusing on limits derived from GW interferometer data.
III Observational constraints on DM directly interacting with GW interferometers
Section II outlined several theoretically motivated models of DM that could leave detectable imprints in GW interferometer data, and we showed how these models map on to induced strains on the detectors. The natural next question is whether current and future interferometers are actually sensitive to such weak signals. In this section, we turn to the search strategies themselves: the methods used to look for DM interactions with gravitational-wave detectors, the observational constraints that have already been obtained, and the prospects for upcoming instruments.
Each method introduced in Section III.1 can be characterized by a sensitivity, i.e. the minimum detectable strain amplitude at a given confidence level. This quantity is what ultimately gets translated into limits on the physical couplings between DM and the standard model, and does not typically depend on the DM model considered. In other words, these search techniques are model-agnostic: they look for correlated, excess power at each frequency the dataset, independent of the underlying nature of DM. The interpretation comes later, when the observed sensitivities are framed as constraints on particular coupling constants and interaction scenarios.
After describing the methods, we present current observational constraints set by ground- and space-based instruments in Section III.2, and those set by pulsar timing arrays in Section III.3. We conclude in Section III.4 by discussing how future GW interferometers could set tighter constraints on, or potentially detect, a wider variety of DM models, as outlined in Section II.
@@ III.1 Methods
If the signal were purely monochromatic, we could simply take a single fast Fourier transform of the data and look for peaks in the power spectrum. However, the stochastic frequency variations prevent us from doing that. If we observe for a duration $T_{\text{obs}}$ longer than $T_{\text{coh}}$, the signal will not be sinusoidal and thus its power will be spread among different frequency bins, as shown in Fig. 3, which would inhibit a possible detection. Thus, the following methods have been designed to combine signal power across chunks of data of length $T_{\text{FFT}}\ll T_{\text{obs}}$ in such a way that avoid power loss and thus optimizes sensitivity towards particle DM interactions with the standard model and gravity. Essentially, these methods vary the length of the chunk of data $T_{\text{FFT}}$ to match the coherence time of DM $T_{\text{coh}}$.
Each of the methods described below formulates its sensitivity, detection statistic, and search strategy differently. As such, applying multiple methods to search for the same type of DM is worthwhile, since each independently defines how to assess the significance of candidates, compute upper limits, set detection thresholds, and veto outliers that historically have arisen from instrumental or environmental noise. Using diverse approaches provides a valuable cross-check and increases the robustness of any potential detection or constraint.
@@ III.1.1 Cross-correlation
Conceptually, cross-correlation Pierce et al. (2018); Guo et al. (2019) requires at least two separate time-domain datasets that are Fourier transformed and multiplied together to compute the cross-power in each frequency bin. The cross power is then divided by the auto-power (the power spectral density) of each detector, and then summed over all the Fourier transforms to arrive at a measure of power at each frequency. It is then divided by the standard deviation of the noise to compute the signal-to-noise ratio (SNR).
Mathematically, the cross-correlated signal strength for detector pair $IJ$ is Pierce et al. (2018); Guo et al. (2019):
@@
in the $j^{\rm th}$ frequency bin at the $i^{\rm th}$ time. “*” denotes the complex conjugate, and $z_{I,ij}$ and $z_{J,ij}$ is the Fourier transform of the time-domain data from detectors II and $J$, respectively, and $N_{\text{FFT}}$ is the total number of FFTs taken over the observing run. If the observing run lasts $T_{\text{obs}}$ and the FFT length is $T_{\text{FFT}}$, then $N_{\text{FFT}}=T_{\text{obs}}/T_{\text{FFT}}$. $P_{I,ij}$ and $P_{J,ij}$ are the individual detector power spectral densities, which are typically estimated using a running median over each FFT. Typically, $T_{\text{FFT}}=1800$ s, independently of the DM mass, which ensures a relatively straightforward analysis but implies the sensitivity towards DM interactions is not optimal at each mass analyzed.
In addition to the signal strength, we have to obtain an estimate of the variance of the noise data. This can be computed from the individual power spectral densities in each FFT as Guo et al. (2019):
@@
The SNR, the detection statistic in each frequency bin, is then simply:
@@
In the presence of pure Gaussian noise, the SNR will follow a normal distribution, with a mean of 0 and a standard deviation of 1. If the SNR exceeds a certain threshold, which is set both theoretically assuming Gaussian noise and by the trials factor (accounting for the size of the parameter space), then a particular frequency is classified as being “significant”.
Depending on the type of ultralight dark matter being searched for, the detectability of a signal in cross-correlation searches between interferometers depends on the spin of the DM: spin-0 (scalar), spin-1 (vector), or spin-2 (tensor) Manita et al. (2024). The SNR for a given detector pair is determined by the DM energy density fraction in that spin state, the strength of its couplings to standard-model particles, the observation time, and the relative orientation, location and geometry of the detectors. The latter is characterized by an “overlap reduction function” (ORF), which determines the sensitivity loss when applying cross-correlation to non-collocated, non-aligned detector pairs. For these models of DM, two kinds of ORFs arise: one from the spatial displacement of mirrors, and another from the finite-light time travel effects of the signal on the mirrors, a previously explained in Section II.4 Morisaki et al. (2021). In particular, cross-correlation is not nearly as sensitive to the finite-light travel time effect as to the spatial effect when it was applied to search for spin-1 DM, which will be shown in Section III.2.
Larger couplings, longer effective observation times, and favorable overlap between detectors increase the SNR and thus the probability of detecting a signal. Conversely, weaker couplings, shorter observations, or misaligned detectors reduce the SNR, making a signal more difficult to distinguish from instrumental noise. After an SNR is computed via Eq. 68, it can be translated directly from SNR to coupling strengths using Eq. 120-122 in Manita et al. (2024)).
It is also useful to calculate a minimum detectable strain amplitude, which can also be used to place constraints on the DM coupling constants for scalar, vector, or tensor ultralight dark matter. Given the SNR, the corresponding strain amplitude can be estimated as:
$h_{0,j}=\sqrt{\frac{2\,\mathrm{SNR}\,P_{I,j}\,P_{J,j}}{\gamma^{2}\,T_{\text{obs}}\,T_{\text{FFT}}}},$ (68)
where $\gamma$ is the ORF that depends on the way that different kinds of DM couple to the interferometers Manita et al. (2024). By inserting the SNR returned from a search (Eq. 67), we can obtain an estimate of the minimum detectable strain amplitude for a potential DM signal, and then map those amplitude values to constraints on DM couplings in spin-0, spin-1 or spin-2 using the strain amplitude equations derived in Section II, analogously to translating the SNR directly to coupling strengths in Manita et al. (2024).
III.1.2 Excess power method